# Rateless Stochastic Coding for Delay-constrained Semantic Communication

### v1

@@ Abstract
We consider the problem of joint source-channel coding with distortion and perception constraints from a rateless perspective, the purpose of which is to settle the balance between reliability (distortion/perception) and effectiveness (rate) of transmission over uncertain channels. We find a new finite-blocklength bound for the achievable joint source-channel code rate with the above two constraints. To achieve a superior rateless characteristic of JSCC coding, we perform multi-level optimization on various finite-blocklength codes. Based on these two, we then propose a new JSCC coding scheme called rateless stochastic coding (RSC). We experimentally demonstrate that the proposed RSC can achieve variable rates of transmission maintaining an excellent trade-off between distortion and perception.
@@ I Introduction
It is well known that the trade-off between distortion, perceptual quality, and coding rate is generally a problem of source compression. Considering that in rateless codes the codeword assigned to each message is always infinite, while the actual effective length depends on the channel conditions and the receiver’s decoding strategy. Based on the above observations, we believe that trying to apply the rateless code directly to an uncompressed source to achieve good compression and reliable communication simultaneously is a meaningful problem, Therefore, we consider the rateless strategy in Joint Source-Channel Coding(JSCC). Furthermore, to satisfy the delay constraints, we analyze the bounds on the distortion and the perceptual quality that can be optimized in the time-limited decoding, which provides a theoretical basis for our actual coding program.
@@
  • The characterization of JSCC with distortion and perception constraints in the finite blocklength regime: We first define $(\epsilon,\gamma,n)$-achievability that is related to distortion, perception, and coding blocklength, then show the achievable capacity region of JSCC in the finite blocklength regime.
  • New rateless stochastic coding scheme: We propose a novel JSCC scheme for semantic communication. This coding scheme considers both the symbol-level information and the global semantic statistical characteristics in the message. To enhance semantic perception, we introduce common randomness to implement stochastic encoding and decoding. Additionally, through the characteristics of rateless codes, the scheme achieves effective compression and transmission of the above two types of information over various coding rates.
@@ II-B Rateless code
III Problem Formulation and Main Results
@@ III-A System Model
We consider the semantic communication system introduced in [15] as shown in Fig. 2. However, it is important to note that we take into account the limitation of delay in channel transmission. In this system, a semantic information source involves some intrinsic knowledge or states that cannot be directly observed by the encoder. By adopting some sampling techniques, the transmitter can collect a limited number of indirect observations of semantic source, i.e., explicit signals that reflect some properties or attributes of the semantic information source. The transmitter will then compress the indirect observations using a joint source-channel encoder $f$ according to the channel capacity to be sent to the receiver. The main objective is to recover the complete semantic information including all the intrinsic knowledge with minimized semantic dissimilarity, called semantic distance, at the receiver.
More formally, the semantic information is formulated as an $m$-length i.i.d. random variable, $W^{m}$, generated by the semantic information source. The encoder can obtain a limited number $n$ of indirect observations of the semantic information source, denoted by $U^{n}$ The indirect observation of the encoder may correspond to a limited number of random samples of noisy observations about the semantic information source obtained by the encoder. Let $p_{W|U}$ be the conditional probability induced by the indirect observation at the transmitter under a given semantic information source $W$. Due to the limited channel capacity and maximum latency $M$, The signals sent to the channel by the encoder contain only $k\leq M$ valid sequences, denoted as $X^{k}$.
In addition to their input signals, both the encoder and the decoder can also have access to two types of side information:
Figure 1: Illustration of a JSCC problem over a noisy communication channel.
Common Randomness (CR): is a source of randomness that is involved in the encoding function to generate the coded messages sent to the channel, but independent of the semantic source. This may correspond to the stochastic encoder in which the input of the encoder includes both indirect observation and a source of randomness. The source of randomness can be the result of a randomized quantization process included in the encoding process for compressing the indirect observation. Although the source of common randomness is not a costly resource in practical compression scenarios, we characterize the amount of common randomness as a random integer distributed uniformly over $[2^{nR_{c}}]$, which is useful for the theoretical analysis of our main results.
Side Information (SI): corresponds to some background information related to the semantic information source. For example, in the human communication scenario[16], different human users may use different languages or preferred sequences of words to express the same idea. In this case, side information may correspond to the language background and preference of users when converting the semantic information source, i.e., ideas, to a sequence of words. Generally speaking, both encoder and decoder may have some side information. The side information at the encoder and decoder does not have to be the same. Let $S$ be the side information at the encoder and decoder, respectively.
III-B $(\epsilon,\gamma,n)$-definition
Under the maximum delay limit $M$, we must consider the existence of finite-length codes. This implies that traditional asymptotic properties may not hold. To show the tradeoff between the transmission error and the finite blocklength of the recovered sequence in the rateless coding scheme, we introduce the notion of the achievability of semantic communication in the regime of finite blocklength with excess distortion and perception. Firstly, we give the formal definition of $n-$length encoder and decoder as follows:
@@ III-B $(\epsilon,\gamma,n)$-definition
For an arbitrary set $\mathcal{W}^{m}$, a (stochastic) encoder is a function
${e}_{n}:\mathcal{U}^{n}\times\mathcal{S}\times\{1,2,...,2^{nR_{c}}\}\to\mathcal{X}^{k},$ (1)
@@
${g}_{n}:\mathcal{X}^{k}\times\mathcal{S}\times\{1,2,...,2^{nR_{c}}\}\to\hat{\mathcal{W}}^{m}.$ (2)
Definition 2
Given $\epsilon>0$, $\gamma>0$, $n$, a rate-distortion-perception tuple $(R,R_{c},D,P)$ is $(\epsilon,\gamma,n)$-achievable with strong perception constraint if there exists a code $({f}_{n},{g}_{n})$ that satisfies both the distortion constraint
$\mathbb{P}\{d(W^{m},\hat{W}^{m})>D\}\leq\epsilon,$ (3)
and the perception constraint
$d_{TV}(p_{W^{m}},p_{\hat{W}^{m}})\leq P+\gamma.$ (4)
where $d_{TV}$ is the total variation distance.
@@ IV RSC FOR DELAY-CONSTRAINED COMMUNICATION
In this section, we present our RSC scheme, which can realize the distion-perception trade-off, while performing rateless transmission for the delay-constrained transmission. In addition, we introduce the dithered quantization multi-objective optimization, which plays a key role in the stochastic and rateless characteristic of RSC.
@@ IV-A Design of RSC
Our design of RSC is inspired by the recent successful application of deep NNs (DNNs), and autoencoders, in particular, to the problem of distribution-preserving lossy compression [17], as well as by the promising initial results in the design of end-to-end communication systems using autoencoder architectures [18] [19]. Fig.2 displays a structure of RSC in specific applications, e.g., image transmission. With a slight abuse of notation, the encoder maps n-dimensional input image $x$ to a quantized $M$-length feature vector $z_{q}$. Following the encoding operation, the source-channel coded $z_{q}$ will lose the tail information by the tail drop operation, which effectively simulates transmission with a maximum delay of $M$, allowing for decoding at different rates depending on the channel’s state.
@@ IV-B Dithered Quantization
Then dithered quantization, a kind of quantization with additive uniform noise, is applied to the output of the perceptron $y$, i.e., $z_{q}=\left\lfloor\mathbf{y}-\mathbf{u}\right\rceil$, where $\mathbf{u}\sim\mathcal{U}\left([-\frac{\Delta}{2},\frac{\Delta}{2})^{n}\right)$ and ${\Delta}$ is minimum quantization interval, i.e., $\Delta=\frac{2}{L_{q}-1}$. In the experiment, we set the quantization level to $L_{q}=4$ and the dimension of $z_{q}$ is set to 12 on the MINIST dataset. Accordingly, the decoder will add the same noise before decoding with the assumption that $u$ is available to both the encoder and the decoder as a common randomness. Specifically, we generate noise by using a pseudorandom number generator with the same random seed. Without considering the channel noise and coding rate changes, the entire process of noise addition and denoising can be expressed as[20]:
$\left\lfloor\mathbf{y}+\mathbf{u}\right\rceil-\mathbf{u}\sim\mathbf{y}+\mathbf{u}^{\prime},$ (5)
@@
IV-C Multi-objective Optimization
The use of Theorem 1 in a specific rateless implementation requires careful consideration of an appropriate optimization method. The distortion between $w$ and the decoder output reconstructed image $\hat{w}$ can be defined by the MSE as: $d(w,\hat{w})=\|w-\hat{w}\|^{2}$. To aid in the assessment and optimization of perceived quality, we adopt the Wasserstein distance, which is the minimum cost of transforming one distribution into another. Specifically, the Wasserstein distance between the source distribution $p_{W}$ and the reconstruction distribution $p_{\hat{w}}$ can be approximated by calculating the output of the discriminator in Wasserstein generative adversarial networks (WGAN)[21]. Since the optimization goal of the discriminator is to maximize the distance between the two distributions, we evaluate the perceptual quality of our reconstruction by pre-training an optimal discriminator $f_{d}^{\ast}$. Therefore, the perceptual distance in the experiment is defined as follows:
$d_{P}\left(p_{w},p_{\hat{w}}\right)=\mathbb{E}_{\hat{w}\sim P_{\hat{W}}}\left[f_{d}^{\ast}(\hat{w})\right]-\mathbb{E}_{w\sim P_{W}}\left[f_{d}^{\ast}(w)\right].$ (6)
Achieving an appropriate Distortion-Perception trade-off is a key point in the design of the JSCC. We first consider the optimization problem under a fixed coding rate. Based on the encoder input $W$ and decoder output $\hat{W}=f_{g}(f_{e}(W))$, We consider the following optimization problem:
$\min_{f_{e},f_{g}}d_{P}\left(P_{\hat{W}},P_{W}\right)+\lambda\mathbb{E}_{W,f_{g}}[d(W,\hat{W})],$ (7)
where $\lambda$ is a hyperparameter. And the loss function of fixed coding rate ${R}$ can be expressed as:
$\min_{f_{e},f_{g}}\overline{\mathcal{L}}_{R}(f_{e},f_{g})=\min_{f_{e},f_{g}}(\underbrace{\mathbb{E}\left[-\log\left(f_{d}^{\ast}\left(f_{g}(f_{e}(w;R)\right),w\right)\right]}_{\text{perception loss }}\\ +\lambda\underbrace{\mathbb{E}\left[d\left(w,f_{g}(f_{e}(w;R)\right)\right]}_{\text{distortion loss }}).$ (8)
In order to better realize the rateless characteristics of our proposed scheme, given the maximum coding rate M, we need to consider the following multi-objective problem:
$\min_{f_{e},f_{g}}[\overline{\mathcal{L}_{1}}_{(}f_{e},f_{g}),\overline{\mathcal{L}_{2}}(f_{e},f_{g}),\ldots,\overline{\mathcal{L}_{M}}(f_{e},f_{g})].$ (9)
By assigning different weights to each rate optimization, we express the multi-objective optimization as the following single objective loss function:
$\min_{f_{e},f_{g}}\sum_{R=1}^{M}\omega_{R}\overline{\mathcal{L}}_{R}(f_{e},f_{g}),$ (10)
where ${\omega_{R}\geq 0}$. The choice of weight distribution should be selected in accordance with the specific JSCC scheme.
Thus far, based on the construction of finite-length codewords from Theorem 1 and the optimization approach mentioned earlier, we have developed a new rateless stochastic coding scheme (RSC), which is characterized by its adaptability to channel conditions without a fixed rate, as well as the stochasticity due to shared common randomness. In the next part, we will conduct a detailed simulation and analysis of RSC.
@@ V-A Simulation Setup
The distribution of dropped tails used in the experiment is power cumulative distribution, the same as the method in [22]. Given the maximum coding rate $M$, the probability distribution of the actual transmission coding rate $D$ satisfies: $Pr(D<{\psi}M)={\psi}^{\beta}$, where $\psi$ is the ratio of practical rate to maximum rate $M$ and $\beta$ is set to 0.67 in all experiments to demonstrate the performance of our model without generality. Accordingly, we allocate weights to the loss function of each rate by ${\omega_{R}=Pr({D=M-R})}$ in (15). Note that unless otherwise noted, the Rate $R$ in our experiment refers to the dimension of the latent dimension after the dropout operation, and the actual coding rate is $\frac{R\times\log_{2}{L_{q}}}{s}$ bits per pixel (bpp) given the image size $s$. In addition, the coding rate control unit requires the ACK signal passed by the receiver to ensure that the sent rateless code is successfully decoded and the sender sends the next codeword.
In the experiment, we only consider the most widely used additive white Gaussian noise(AWGN) channel in our communication system. Specifically, the symbols encoded by the stochastic encoder will be passed through a fixed additive noise module to simulate channel transmission. This process can be expressed as $\hat{z}=z+n_{c}$, $n_{c}\sim\mathcal{N}(0,\sigma^{2}I_{n})$.
@@ VI Conclusions
In this paper, we first showed the JSCC achievability bound with distortion and perception constraints in the finite blocklength regime, which provides a theoretical basis for rateless JSCC schemes under two metrics. Secondly, we proposed rateless stochastic coding (RSC) based on the aforementioned theory, a novel delay-constrained coding scheme that enables rateless while achieving a favorable distortion-perception trade-off. Thirdly, the effectiveness of the proposed RSC was verified by simulating practical communication systems, showing it combines the benefits of stochastic coding and rateless coding. The RSC allowed receivers to obtain essentially artifact-free reconstructions at all rates, maintaining the distribution of original data from learning a generative decoder at low rates or low SNRs on the one hand, learning a decoding strategy with almost perfect reconstruction while considering distortion and perception at high rates or high SNRs on the other hand. That is, the proposed RSC is robust to rate reduction and noise in the AWGN channel.

### v2

@@ Abstract
We consider the problem of joint source-channel coding for semantic communication from a rateless perspective, the purpose of which is to settle the balance between reliability (distortion/perception) and effectiveness (rate) of transmission over uncertain channels. In particular, we propose a more general communication objective that minimizes the perceptual distance by incorporating a semantic-level reconstruction objective in addition to the conventional pixel-level reconstruction objective. Based on the proposed objective, we then propose a novel JSCC coding scheme called rateless stochastic coding (RSC) by introducing a generative decoder and dithered quantization. The coding scheme enables reconstruction based on both distortion and perception metrics through rateless transmission. Extensive experiments demonstrate that the proposed RSC can achieve variable rates of transmission maintaining an excellent trade-off between distortion and perception.
Index Terms:
Semantic communication, distortion, perception, rateless coding.
@@ I Introduction
  • General optimizing objective of JSCC for semantic communication: We propose a general optimizing objective of JSCC for semantic communication under delay constrains, in which not only minimum distortion is needed to achieve pixel-level reconstruction in the classical scheme, but also semantic-level reconstruction by minimizing perception.
  • Novel rateless stochastic coding scheme: We propose RSC, a novel rateless stochastic coding scheme for minimizing the proposed objective, where MSE was used to measure distortion and Wasserstein distance to measure perceived distance, respectively. In addition, we employ a tail drop operation on the encoded feature in order to achieve rateless coding.
@@ II-B Rateless code
III System Model and Problem Formulation
@@ III-A System Model
We consider the end-to-end semantic communication system introduced in [15] as shown in Fig. 1. However, it can be noted that we take into account the limitation of delay in channel transmission and remove the side information setting.
In this system, the encoder can obtain a limited number $n$ of indirect observations of the semantic information source, denoted by $W^{n}$. The transmitter will then compress the indirect observations to $Z^{k}$ using a joint source-channel encoder $f$. Due to the limited channel capacity and maximum latency $K$, The signals sent to the channel by the encoder contain only $k\leq K$ valid sequences, denoted as $Z^{k}$. After transmission, the joint source-channel decoder $g$ decodes input $\hat{Z}^{k}$ to realize reconstruction $\hat{W}^{n}$.
Figure 1: Illustration of a semantic communication system with CR over a noisy communication channel.
In addition to the input signals, both the encoder and the decoder can access common randomness (CR), which is a source of randomness involved in the encoding and decoding function, but independent of the semantic source. This may correspond to the stochastic encoder in which the encoder’s input includes both indirect observation and a source of randomness. The source of randomness can be the result of a randomized quantization process included in the encoding process for compressing the indirect observation. Although the source of common randomness is not a costly resource in practical compression scenarios, we characterize the amount of common randomness as a random integer distributed uniformly over $[2^{nR_{c}}]$, which is useful for the theoretical analysis. Specifically, the common randomness can be realized by sharing some seeds of the pseduo-random number generator in advance.
III-B Problem Formulation
The formal definition of an encoder and decoder is as follows:
@@ III-B Problem Formulation
For an arbitrary set $\mathcal{W}^{n}$, a (stochastic) encoder is a function
${f}:\mathcal{W}^{n}\times\{1,2,...,2^{nR_{c}}\}\to\mathcal{Z}^{k},$ (1)
@@
${g}:\hat{\mathcal{Z}}^{k}\times\{1,2,...,2^{nR_{c}}\}\to\hat{\mathcal{W}}^{n}.$ (2)
For simplicity, we omit the superscript $m$, and define reconstruction of $W$ under delay $k$ as $\hat{W}_{k}$ Then we can minimize the following objective function given $k$:
$\mathcal{L}_{k}(f,g)=d_{P}\left(P_{W},P_{\hat{W}_{k}}\right)+\lambda\mathbb{E}_{W,g}[d(W,\hat{W}_{k})],$ (3)
where $d_{P}\left(P_{W},P_{\hat{W}_{k}}\right)$ is any perceptual measure and $d(W,\hat{W}_{k})$ is any distortion measure. Note that in our communication goal not only minimizing distortion is needed to achieve pixel-level reconstruction in the classical scheme, but also semantic-level reconstruction by minimizing perception Naturally, given maximum latency $K$, we should minimize:
$\mathcal{L}_{\text{total}}(f,g)=\sum_{k=1}^{K}\omega_{k}\mathcal{L}_{k},$ (4)
where $\omega_{k}\geq 0$.
@@ IV RSC For Delay-constrained Communication
In this section, we present our RSC scheme, which can realize the distion-perception trade-off, while performing rateless transmission for the delay-constrained transmission. In addition, we introduce the dithered quantization and multi-objective optimization, which plays a key role in the stochastic and rateless characteristic of RSC.
@@ IV-A Design of RSC
Our design of RSC is inspired by the recent successful application of deep NNs (DNNs), and autoencoders, in particular, to the problem of distribution-preserving lossy compression [16], as well as by the promising initial results in the design of end-to-end communication systems using autoencoder architectures [17] [18]. Fig.2 displays a structure of RSC in specific applications, e.g., image transmission. With a slight abuse of notation, the encoder $f_{\phi}$ parameterized by $\phi$ maps $m$-dimensional input image $w$ to a quantized feature vector $z_{q}$. Following the encoding operation, $z_{q}$ will lose the tail information by the tail drop operation, which effectively simulates rateless transmission with a delay of $k$, allowing for being decoded at different rates depending on the channel’s state by the decoder $g_{\theta}$ parameterized by $\theta$.
@@ IV-B Dithered Quantization
Then dithered quantization, a kind of quantization with additive uniform noise, is applied to the output of the perceptron $y$, i.e., $z_{q}=\left\lfloor y-u\right\rceil$, where $u\sim\mathcal{U}\left([-\frac{\Delta}{2},\frac{\Delta}{2})^{n}\right)$ and ${\Delta}$ is minimum quantization interval, i.e., $\Delta=\frac{2}{L_{q}-1}$. Accordingly, the decoder will add the same noise before decoding with the assumption that $u$ is available to both the encoder and the decoder as a common randomness. Specifically, we generate noise by using a pseudorandom number generator with the same random seed. Without considering the channel noise and coding rate changes, the entire process of noise addition and denoising can be expressed as[19]:
$\left\lfloor y+u\right\rceil-u\sim y+u^{\prime},$ (5)
@@
IV-C Optimization of RSC
We next discuss how to optimize (3) further. The distortion between $w$ and the reconstructed image $\hat{w}_{k}$ can be defined by the MSE as: $d(w,\hat{w}_{k})=\|w-\hat{w}_{k}\|^{2}$. To aid in the assessment and optimization of perceived quality, we adopt the Wasserstein distance, which is the minimum cost of transforming one distribution into another. Specifically, the Wasserstein distance between the source distribution $p_{w}$ and the reconstruction distribution $p_{\hat{w}_{k}}$ can be approximated by calculating the output of the discriminator in Wasserstein generative adversarial networks (WGAN)[20]. Since the optimization goal of the discriminator is to maximize the distance between the two distributions, we evaluate the perceptual quality of our reconstruction by pre-training an optimal discriminator $r_{\zeta}^{\ast}$ parameterized by $\zeta$. Therefore, the perceptual distance in the experiment is defined as follows:
$d_{P}\!\left(p({\hat{w}_{k}}),p({w})\right)\!=\!\mathbb{E}_{\hat{w}_{k}\sim p({\hat{w}_{k}})}\!\left[r_{\zeta}^{\ast}(\hat{w}_{k})\right]\!-\!\mathbb{E}_{w\sim p(w)}\!\left[r_{\zeta}^{\ast}(w)\right].$ (6)
In the proposed RSC, we consider the most widely used additive white Gaussian noise(AWGN) channel in our communication system. Specifically, the symbols encoded by the stochastic encoder will be passed through a fixed additive noise module to simulate channel transmission. This process can be expressed as $\hat{z}=z+n_{c}$, $n_{c}\sim\mathcal{N}(0,\sigma^{2}I_{n})$. Then we can rewrite (3) as:
${\mathcal{L}}_{k}(f_{\phi},g_{\theta})=\underbrace{\mathbb{E}\left[-\log\left(r_{\zeta}^{\ast}\left(g_{\theta}(f_{\phi}(w;k)+n_{c}\right),w\right)\right]}_{\text{perception loss }}\\ +\lambda\underbrace{\mathbb{E}\left[d\left(w,g_{\theta}(f_{\phi}(w;k)+n_{c})\right)\right]}_{\text{distortion loss }}.$ (7)
Accordingly, minimizing (4) can be rewritten as:
${\mathcal{L}_{\text{total}}}(f_{\phi},g_{\theta})\!=\!\sum_{k=1}^{K}\omega_{k}\bigg{\{}{\!\mathbb{E}\!\left[\!-\!\log\left(s_{\zeta}^{\ast}\left(g_{\theta}(f_{\phi}(w;k)\!+\!n_{c}\right),w\right)\right]}\\ +\lambda{\mathbb{E}\left[d\left(w,g_{\theta}(f_{\phi}(w;k)+n_{c})\right)\right]}\bigg{\}}.$ (8)
Thus far, we have developed a new rateless stochastic coding scheme (RSC), which is characterized by its adaptability to channel conditions without a fixed rate, as well as the stochasticity due to shared common randomness. In the next part, we will conduct a detailed simulation and analysis of RSC.
@@ V-A Simulation Setup
The distribution of dropped tails used in the experiment is power cumulative distribution, the same as the method in [21]. Given the maximum coding rate $K$ and any rate $k$, the probability distribution of the actual drop rate $D=K-k$ satisfies: $Pr(D<{\psi}K)={\psi}^{\beta}$, where $\psi$ is the ratio of practical rate to maximum rate $K$ and $\beta$ is set to 0.67 in all experiments to demonstrate the performance of our model without generality. Accordingly, we allocate weights to the loss function of each rate by ${\omega_{k}=Pr({D})}$ in (4) and (8). Note that unless otherwise noted, the Rate $K$ in our experiment refers to the dimension of the latent dimension after the dropout operation, and the actual coding rate is $\frac{R\times\log_{2}{L_{q}}}{s}$ bits per pixel (bpp) given the image size $s$. In addition, the coding rate control unit requires the ACK signal passed by the receiver to ensure that the sent rateless code is successfully decoded and the sender sends the next codeword.
@@ VI Conclusions
In this paper, we consider a semantic communication system with CR. In contrast to conventional methodologies that solely prioritizes the minimization of distortion, we propose a new communication objective which entails the incorporation of a perceptual component into the objective, thereby facilitating the reconstruction of semantic information at the semantic level. Then we propose rateless stochastic coding (RSC) based on the aforementioned theory, a novel delay-constrained coding scheme that enables rateless while achieving a favorable distortion-perception trade-off. At last, the effectiveness of the proposed RSC was verified by simulating practical communication systems, showing it combines the benefits of stochastic coding and rateless coding. The RSC allowed receivers to obtain essentially artifact-free reconstructions at all rates, maintaining the distribution of original data from learning a generative decoder at low rates or low SNRs on the one hand, learning a decoding strategy with almost perfect reconstruction while considering distortion and perception at high rates or high SNRs on the other hand. That is, the proposed RSC is robust to rate reduction and noise in the AWGN channel.